### v1

# Convergence of empirical Gromov-Wasserstein distance

Abstract.
We study rates of convergence for estimation of the Gromov-Wasserstein distance. For two marginals supported on compact subsets of $\mathbb{R}^{d_{x}}$ and $\mathbb{R}^{d_{y}}$, respectively, with $\min\{d_{x},d_{y}\}>4$, prior work established the rate $n^{-\frac{2}{\min\{d_{x},d_{y}\}}}$ for the plug-in empirical estimator based on $n$ i.i.d. samples. We extend this fundamental result to marginals with unbounded supports, assuming only finite polynomial moments. Our proof techniques for the upper bounds can be adapted to obtain sample complexity results for penalized Wasserstein alignment that encompasses the Gromov-Wasserstein distance and Wasserstein Procrustes in unbounded settings. Furthermore, we establish matching minimax lower bounds (up to logarithmic factors) for estimating the Gromov-Wasserstein distance.
Key words and phrases:
Gromov-Wasserstein distance, minimax lower bound, optimal transport, sample complexity, unbounded support
K. Kato is partially supported by NSF grant DMS-2413405.
1. Introduction
1.1. Overview
The Gromov-Wasserstein (GW) distance [Mém11, Stu12] provides a powerful tool for comparing and aligning heterogeneous and structured data sets and has received increasing interest from various application domains. Examples of applications include shape and graph matching [Mém09, XLZD19, XLC19] and language alignment [AMJ18]. Generally, the GW distance defines a metric on a space of Polish metric measure spaces modulo measure-preserving isometries [Stu12]. For two Euclidean metric measure spaces $(\mathbb{R}^{d_{x}},\|\cdot\|,\mu)$ and $(\mathbb{R}^{d_{y}},\|\cdot\|,\nu)$ endowed with Borel probability measures $\mu$ and $\nu$, the $(p,q)$-GW distance $\mathsf{GW}_{p,q}(\mu,\nu)$ with $p,q\in[1,\infty)$ is defined by
$\begin{split}\mathsf{GW}_{p,q}^{p}(\mu,\nu)&:=\mathsf{GW}_{p,q}^{p}\Big{(}(\mathbb{R}^{d_{x}},\|\cdot\|,\mu),(\mathbb{R}^{d_{y}},\|\cdot\|,\nu)\Big{)}\\ &:=\inf_{\pi\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}}\int_{\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}}\big{|}\|x-x^{\prime}\|^{q}-\|y-y^{\prime}\|^{q}\big{|}^{p}\,d\pi(x,y)d\pi(x^{\prime},y^{\prime}),\end{split}$ (1)
where $\Pi(\mu,\nu)$ denotes the set of couplings for $\mu$ and $\nu$.
Despite its widespread applications, the statistical analysis of GW remains challenging. In contrast to classical optimal transport (OT), for which a rich statistical theory exists, GW presents significant obstacles due to the bilinear and nonconvex nature of its objective function, as opposed to linear OT. Additionally, there is still a lack of a comprehensive duality theory for general GW that limits the development of detailed statistical theory.
The recent work [ZGMS24] established the first sample complexity results for the GW distance with $(p,q)=(2,2)$ by leveraging a variational representation of $\mathsf{GW}_{2,2}^{2}$ that links the GW problem to standard OT. Using this representation and OT duality theory, they showed that for compactly supported $\mu$ and $\nu$, the empirical estimator based on $n$ i.i.d. samples converges at the rate $n^{-2/(d_{x}\wedge d_{y})}$ with $d_{x}\wedge d_{y}=\min\{d_{x},d_{y}\}$ when $d_{X}\wedge d_{y}>4$. [^1] The empirical estimator is defined by plugging in the empirical distributions for $\mu$ and $\nu$. Notably, this rate adapts automatically to the smaller of the two dimensions, rather than being governed by the worst case dimension $d_{x}\vee d_{y}=\max\{d_{x},d_{y}\}$. This is achieved by adapting the lower complexity adaptation (LCA) principle studied in [HSM24] in the OT case. They further derived matching lower bounds for the empirical estimator, but did not derive minimax lower bounds. As such, strictly speaking, their lower bound result does not rule out the possibility that estimators other than the plug-in type could uniformly outperform the empirical estimator.
The first goal of this work is to extend their fundamental sample complexity result to the unbounded support case. Our result establishes that, when $d_{x}\wedge d_{y}>4$, the $n^{-2/(d_{x}\wedge d_{y})}$ rate (without logarithmic factors) continues to hold for the empirical GW even when $\mu,\nu$ only possess finite polynomial moments. It is worth noting that, even in OT cost estimation, extending results from the compact support case to the unbounded setting is often highly nontrivial. This is because global regularity estimates for dual potentials, which are often available for the compact support case, do not continue to hold for the unbounded setting, and establishing local regularity estimates would require delicate tail conditions on the marginals; see [CF21, MBNWW24] and the discussion in Section 2. Our proof essentially builds on an adaptation of an idea in [SH25], but with some nontrivial twists; see the discussion after Theorem 3.1 below for details. Furthermore, our proof technique can be adapted to obtain sample complexity results for penalized Wasserstein alignment [PSW25] that encompasses the Gromov-Wasserstein distance and Wasserstein Procrustes in unbounded settings.
Our second goal is to formally derive minimax lower bounds for estimating $\mathsf{GW}_{2,2}^{2}$ that match the upper bounds for the empirical estimator, possibly up to logarithmic factors. For $d_{x}\wedge d_{y}>4$, our result establishes a minimax lower bound that matches $n^{-2/(d_{x}\wedge d_{y})}$ up to a logarithmic factor, for the class of distributions supported in the unit ball (and hence any larger distribution class). This result indicates that no other estimators can significantly outperform the empirical estimator uniformly over the said class of distributions. The proof builds on [NWR22] but requires some new ideas to deal with invariance of GW under isometries.
1.2. Literature review
The literature related to this paper is broad. We refer the reader to [CNWR24] as an excellent monograph on recent developments of statistical OT.
Convergence and exact asymptotics of empirical OT costs have been extensively studied in the statistics and probability literature; see, e.g., [Dud69, AKT84, Tal92, Tal94, DY95, BdMM02, BB13, DSS13, BLG14, FG15, WB19, Lei20, CRL+20, MNW24, SH25]. Most of these references focus on establishing sharp rates for empirical distributions under $p$-Wasserstein distances $W_{p}$ with $p\in[1,\infty)$. [^2] See Section 2 for the definition of $W_{p}$. For instance, let $\mu$ be a Borel probability measure on $\mathbb{R}^{d}$ and $\hat{\mu}_{n}$ be the empirical distribution for $n$ i.i.d. samples from $\mu$; then, [FG15] showed that when $\mu$ has a finite $q$-th moment with $q>p$,
[…]
where $\lesssim$ denotes an inequality holding up to a numerical constant that is independent of $n$ but may depend on other parameters. As a canonical case, when $d>4$ and q>2d/(d2)q>2d/(d-2), the result (2) implies $\mathbb{E}\big{[}W_{2}^{2}(\hat{\mu}_{n},\mu)\big{]}\lesssim n^{-2/d}$. These rates are known to be sharp in various settings with a notable exception of the $d=2p$ case; cf. the discussion after Theorem 1 in [FG15].
Estimation of the OT cost $W_{p}^{p}$, rather than distribution estimation under $W_{p}$, has been explored by the recent works by [CRL+20, MNW24, SH25]. For instance, let $\nu$ be another Borel probability measure on $\mathbb{R}^{d}$ and $\hat{\nu}_{m}$ be the empirical distribution of $m$ i.i.d. samples from $\mu$ that are independent of the samples from $\mu$; then, [CRL+20] showed that, when $d>4$ and $\mu,\nu$ are compactly supported,
$\mathbb{E}\big{[}\big{|}W_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{m})-W_{2}^{2}(\mu,\nu)\big{|}\big{]}\lesssim(n\wedge m)^{-2/d}.$ (3)
The same rate holds for estimating $W_{2}(\mu,\nu)$ as long as $W_{2}(\mu,\nu)$ is bounded away from zero, which is faster than the rate implied by (2) combined with the triangle inequality. See [MNW24, SH25] for extensions to general $p$ and marginals with unbounded supports. Furthermore, the rate in (3) agrees with a minimax lower bound up to a logarithmic factor for a class of distributions supported in a fixed ball [MNW24].
In contrast to standard OT costs, statistical analysis of GW distances is still in its infancy. In addition to GW itself, [ZGMS24] studied entropic regularization of GW with $(p,q)=(2,2)$, establishing parametric sample complexity results analogous to those in entropic OT estimation (cf. [GCB+19, MNW19]). [GH24] studied the LCA principle for entropic GW, focusing on how the power of the regularization parameter depends on the intrinsic dimensionality. A recent preprint [RGK24b] derived the first limiting distributional results for empirical GW in both discrete and semi-discrete settings. The present paper contributes to the (ever-growing) statistical OT literature by deepening the understanding of the fundamental statistical properties of GW, which remain underdeveloped despite significant interest from applied domains.
1.3. Organization
The rest of the paper is organized as follows. Section 2 collects brief overviews of OT and GW and a discussion on the prior sample complexity results for GW. Section 3 presents the main results on sample complexity in the unbounded setting and minimax lower bounds for GW. All proofs are gathered in Section 4.
1.4. Notation
For two numbers $a,b\in\mathbb{R}$, we use the notation $a\wedge b=\min\{a,b\}$ and $a\vee b=\max\{a,b\}$. Let $\|\cdot\|_{\operatorname{op}}$ and $\|\cdot\|_{\mathrm{F}}$ denote the operator and Frobenius norms for matrices, respectively, i.e., for any matrix $A=(a_{ij})_{\begin{subarray}{c}1\leq i\leq d_{1}\\ 1\leq j\leq d_{2}\end{subarray}}\in\mathbb{R}^{d_{1}\times d_{2}}$,
$\|A\|_{\operatorname{op}}:=\sup_{y\in\mathbb{R}^{d_{2}},y\neq 0}\frac{\|Ay\|}{\|y\|}\quad\text{and}\quad\|A\|_{\mathrm{F}}:=\sqrt{\sum_{\begin{subarray}{c}1\leq i\leq d_{1}\\ 1\leq j\leq d_{2}\end{subarray}}a_{ij}^{2}}.$
For any symmetric matrix $A$, let $\lambda_{\min}(A)$ denote its smallest eigenvalue. For any metric space $(M,d)$, we use $B_{M}(x,r)$ to denote the closed ball in $M$ with center $x$ and radius $r$. Let $\mathcal{P}(M)$ denote the collection of all Borel probability measures on $M$. For any $p>0$ and any fixed $x_{0}\in M$, let $\mathcal{P}_{p}(M):=\{\mu\in\mathcal{P}(M):\int d^{p}(x,x_{0})\,d\mu(x)<\infty\}$. For any $\mu\in\mathcal{P}(M)$ and $p\in[1,\infty)$, let $(L^{p}(\mu),\|\cdot\|_{L^{p}(\mu)})$ denote the $L^{p}$-space of Borel measurable real functions on $M$ with respect to (w.r.t.) $\mu$. For any $\mu\in\mathcal{P}(M)$ and any Borel measurable mapping $f$ from $M$ into another metric space, let $f_{\#}\mu$ denote the pushforward of $\mu$ under $f$, i.e., $f_{\#}\mu=\mu\circ f^{-1}$. For real functions $f,g$ defined on spaces $\mathcal{X},\mathcal{Y}$, respectively, let $f\oplus g$ denote their tensor sum, i.e., $(f\oplus g)(x,y)=f(x)+g(y)$. For two probability measures $\mu,\nu$, let $\mu\otimes\nu$ denote their product measure. Finally, the notation $\lesssim$ signifies an inequality that holds up to a numerical constant independent of $(n,m)$ but that may depend on other parameters. The dependence of the hidden constant on the parameters will be clarified from place to place.
2. Preliminaries
Throughout the paper, let $d_{x},d_{y}\in\mathbb{N}$ be fixed. For notational convenience, let $\mathcal{X}=\mathbb{R}^{d_{x}}$ and $\mathcal{Y}=\mathbb{R}^{d_{y}}$. Let $\mu\in\mathcal{P}(\mathcal{X})$ and $\mu\in\mathcal{P}(\mathcal{Y})$ be given. Suppose that there are i.i.d. samples $X_{1},\dots,X_{n}$ and $Y_{1},\dots,Y_{m}$ from $\mu$ and $\nu$, respectively, that are independent of each other. The corresponding empirical distributions are defined by
$\hat{\mu}_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}\quad\text{and}\quad\hat{\nu}_{m}:=\frac{1}{m}\sum_{j=1}^{m}\delta_{Y_{j}}.$ (4)
These notations will be carried over to the next sections. Furthermore, we assume $n\wedge m\geq 2$ (this is to avoid $\log(n\wedge m)=0$).
In this section, we first review OT and GW and move on to discussing prior related results.
2.1. Optimal transport
Let $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$ be a continuous, not necessarily nonnegative, cost function. Assume that there exist nonnegative continuous functions $c_{\mathcal{X}}:\mathcal{X}\to\mathbb{R}_{+}$ and $c_{\mathcal{Y}}:\mathcal{Y}\to\mathbb{R}_{+}$ such that
$|c|\leq c_{\mathcal{X}}\oplus c_{\mathcal{Y}}\quad\text{on \ $\mathcal{X}\times\mathcal{Y}$}.$
Assume that $c_{\mathcal{X}}\in L^{1}(\mu)$ and $c_{\mathcal{Y}}\in L^{1}(\nu)$. The OT cost between $\mu$ and $\nu$ is defined by
$T_{c}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\int c\,d\pi.$ (5)
The preceding moment condition ensures that $T_{c}(\mu,\nu)$ is finite. Furthermore, the following strong duality holds:
$T_{c}(\mu,\nu)=\sup_{\begin{subarray}{c}f\in L^{1}(\mu),g\in L^{1}(\nu)\\ f\oplus g\leq c\end{subarray}}\int f\,d\mu+\int g\,d\nu,$
where the supremum on the right-hand side is attained. See Theorem 5.9 in [Vil08] or Theorem 6.1.5 in [AGS08]. We call any functions $(f,g)$ achieving the supremum above dual potentials. It is worth noting that dual potentials $(f,g)$ can be chosen to be $c$-concave and one of the potentials can be replaced with the $c$-transform of the other. Recall that, for a function $f:\mathcal{X}\to[-\infty,\infty)$ that is not identically -$\infty$, its $c$-transform is defined by $f^{c}(y):=\inf_{x\in\mathcal{X}}\{c(x,y)-f(x)\}$ for $y\in\mathcal{Y}$, and $f$ is called $c$-concave if it agrees with the $c$-transform of some function on $\mathcal{Y}$. Under the current assumption, there exists a $c$-concave function $f\in L^{1}(\mu)$ with $f^{c}\in L^{1}(\nu)$ such that $(f,f^{c})$ are dual potentials.
When $\mathcal{X}=\mathcal{Y}$, the $p$-Wasserstein distance $W_{p}(\mu,\nu)$ with $p\in[1,\infty)$ corresponds to $T_{c}^{1/p}(\mu,\nu)$ with $c(x,y)=\|x-y\|^{p}$, i.e.,
$W_{p}(\mu,\nu):=\inf_{\pi\in\Pi(\mu,\nu)}\left(\int\|x-y\|^{p}\,d\pi(x,y)\right)^{1/p}.$
The $W_{p}$ defines a metric on $\mathcal{P}_{p}(\mathcal{X})$ and metrizes weak convergence plus convergence of $p$-th moments. We refer the reader to [Vil08, AGS08] as excellent references on OT and Wasserstein distances.
2.2. Gromov-Wasserstein distance
For $p,q\in[1,\infty)$, the $(p,q)$-GW distance is defined by (1), where we assume $\mu\in\mathcal{P}_{pq}(\mathcal{X})$ and $\nu\in\mathcal{P}_{pq}(\mathcal{Y})$ to ensure finiteness of $\mathsf{GW}_{p,q}(\mu,\nu)$. If one views $\mathsf{GW}_{p,q}$ as comparing two metric measure spaces $(\mathcal{X},\|\cdot\|,\mu)$ and $(\mathcal{Y},\|\cdot\|,\nu)$, then it is symmetric and satisfies the triangle inequality. Finally, $\mathsf{GW}_{p,q}(\mu,\nu)=0$ if and only if there exists an isometry $f:\mathrm{spt}(\mu)\to\mathrm{spt}(\nu)$ such that $\nu=f_{\#}\mu$ ($\mathrm{spt}(\mu)$ denotes the support of $\mu$). We record a few properties of $\mathsf{GW}_{p,q}$ that will be used later. Set
$m_{q}(\mu):=\int\|x\|^{q}\,d\mu(x)$
for $q>0$.
Lemma 2.1.
Let $p,q\in[1,\infty)$ be arbitrary. The following holds.
  1. (i)
    If the diameter of the support of each of $\mu$ and $\nu$ is at most $L$ for some constant $L>0$, then
    $\mathsf{GW}_{p,q}(\mu,\nu)\leq qL^{q-1}\mathsf{GW}_{p,1}(\mu,\nu).$
  2. (ii)
    Suppose $d_{x}=d_{y}=d$. For any $\mu,\nu\in\mathcal{P}_{pq}(\mathbb{R}^{d})$, we have
    $\mathsf{GW}_{p,q}(\mu,\nu)\leq q\,2^{q+1-\frac{1}{p}+\frac{1}{pq}}\big{(}m_{pq}(\mu)+m_{pq}(\nu)\big{)}^{\frac{q-1}{pq}}W_{pq}(\mu,\nu).$
    Furthermore, for $p=q=2$, if $\mu$ and $\nu$ have covariance matrices $\Sigma_{\mu}$ and $\Sigma_{\nu}$ with smallest eigenvalues $\lambda_{\mathrm{min}}(\Sigma_{\mu})$ and $\lambda_{\mathrm{min}}(\Sigma_{\nu})$, respectively, then
    $\Big{(}32\big{(}\lambda^{2}_{\mathrm{min}}(\Sigma_{\mu})+\lambda^{2}_{\mathrm{min}}(\Sigma_{\nu})\big{)}\Big{)}^{1/4}\inf_{U\in E(d)}W_{2}(\mu,U_{\#}\nu)\leq\mathsf{GW}_{2,2}(\mu,\nu),$
    where $E(d)$ denotes the isometry group on $\mathbb{R}^{d}$.
Part (i) is due to Lemma 9.5 (iii) in [Stu12] and follows directly from the elementary inequality $|a^{q}-b^{q}|\leq qL^{q-1}|a-b|$ for $a,b\in[0,L]$. Part (ii) is due to Lemma 4.4 in [ZGMS24]. [^3] Lemma 4.4 in [ZGMS24] assumes that $\Sigma_{\mu}$ and $\Sigma_{\nu}$ are of full rank, but this assumption can be removed.
In this paper, as in [ZGMS24, RGK24a, GH24], we focus on the $(p,q)=(2,2)$ case. For notational convenience, set
$D(\mu,\nu):=\mathsf{GW}_{2,2}^{2}(\mu,\nu).$
The following notation will be useful:
$\begin{split}S_{1}(\mu,\nu)&:=\int\|x-x^{\prime}\|^{4}\,d\mu\otimes\mu(x,x^{\prime})+\int\|y-y^{\prime}\|^{4}\,d\nu\otimes\nu(y,y^{\prime})\\ &\qquad-4\int\|x\|^{2}\|y\|^{2}\,d\mu\otimes\nu(x,y),\\ S_{2}(\mu,\nu)&:=\inf_{\pi\in\Pi(\mu,\nu)}\bigg{\{}\int-4\|x\|^{2}\|y\|^{2}\,d\pi(x,y)-8\left\|\int xy^{\top}\,d\pi(x,y)\right\|_{\mathrm{F}}^{2}\bigg{\}}.\end{split}$
Expanding the squares, one can decompose $D(\mu,\nu)$ as
$D(\mu,\nu)=S_{1}(\bar{\mu},\bar{\nu})+S_{2}(\bar{\mu},\bar{\nu}),$
where $\bar{\mu}$ and $\bar{\nu}$ are the centered versions of $\mu$ and $\nu$, respectively, i.e., $\bar{\mu}$ is the distribution of $X-\mathbb{E}[X]$ when $X\sim\mu$. The first term, $S_{1}$, only involves moments of the marginals and is not difficult to handle. For the analysis of the second term $S_{2}$, the following variational representation, due to [ZGMS24], is particularly useful, as it allows us to link the GW problem to standard OT. A variant of the variational representation also plays a key role in developing formal computational guarantees for entropic GW; see [RGK24a].
Lemma 2.2 (Variational representation; Corollary 4.1 in [ZGMS24]).
For any $\mu\in\mathcal{P}_{4}(\mathcal{X})$ and $\nu\in\mathcal{P}_{4}(\mathcal{Y})$, one has
$S_{2}(\mu,\nu)=\inf_{A\in\mathbb{R}^{d_{x}\times d_{y}}}\big{\{}32\|A\|_{\mathrm{F}}^{2}+T_{c_{A}}(\mu,\nu)\big{\}},$ (6)
where $c_{A}(x,y):=-4\|x\|^{2}\|y\|^{2}-32x^{\top}Ay$. Furthermore, the infimum on the right-hand side is achieved by some $A\in\mathbb{R}^{d_{x}\times d_{y}}$ with $\|A\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}/2$.
The proof is simple and based on the observation that
$8\left\|\int xy^{\top}\,d\pi(x,y)\right\|_{\mathrm{F}}^{2}\leq 32\|A\|_{\mathrm{F}}^{2}-32\left\langle A,\int xy^{\top}\,d\pi(x,y)\right\rangle_{\mathrm{F}}$
with equality holding if and only if $A=\frac{1}{2}\int xy^{\top}\,d\pi(x,y)$, where $\langle\cdot,\cdot\rangle_{\mathrm{F}}$ denotes the Frobenius inner product and $\langle A,\int xy^{\top}\,d\pi(x,y)\rangle_{\mathrm{F}}=\int x^{\top}Ay\,d\pi(x,y)$. Interchanging $\inf_{\pi}$ and $\inf_{A}$ gives the expression (6). Finally, since $\|\int xy^{\top}\,d\pi(x,y)\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}$, $\inf_{A}$ in (6) can be reduced to the infimum over $A$ with $\|A\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}/2$. As the mapping $A\mapsto 32\|A\|_{\mathrm{F}}^{2}+T_{c_{A}}(\mu,\nu)$ is continuous, the final claim follows.
2.3. Prior results for upper bounds and challenges in unbounded settings
A fundamental statistical question is estimation of $D(\mu,\nu)$ from samples. A natural estimator is the empirical estimator $D(\hat{\mu}_{n},\hat{\nu}_{m})$. Theorem 4.2 in [ZGMS24] establishes the following sample complexity bound when $\mu$ and $\nu$ are supported in $B_{\mathcal{X}}(0,r)$ and $B_{\mathcal{Y}}(0,r)$, respectively, for some $r\geq 1$:
$\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim r^{4}\varphi_{n,m},$ (7)
where
$\varphi_{n,m}:=(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}(\log(n\vee m))^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}.$ (8)
The hidden constant in (7) depends only on $d_{x}$ and $d_{y}$. Precisely speaking, [ZGMS24] only considered the $n=m$ case but the $n\neq m$ case follows similarly with a minor modification.
The proof of Theorem 4.2 in [ZGMS24] leverages the variational representation from Lemma 2.2 and OT duality theory. Given the variational representation, the approach is similar to the one used in [CRL+20]. In the GW case, exploiting the variational representation, [ZGMS24] reduces the problem to bounding
$\sup_{A}\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|},$ (9)
where $\sup_{A}$ is taken over a compact subset of $\mathbb{R}^{d_{x}\times d_{y}}$. Suppose that $4<d_{x}\leq d_{y}$, so that $d_{x}\wedge d_{y}=d_{x}$. Using OT duality, [ZGMS24] further reduces the problem to finding upper bounds on
$\sup_{f\in\mathcal{F}}\left|\int f\,d(\hat{\mu}_{n}-\mu)\right|\quad\text{and}\quad\sup_{g\in\mathcal{G}}\left|\int g\,d(\hat{\nu}_{m}-\nu)\right|.$ (10)
where $\mathcal{F}$ and $\mathcal{G}$ are function classes chosen to contain dual potentials for $(\hat{\mu}_{n},\hat{\nu}_{m})$ and $(\mu,\nu)$ w.r.t. cost $c_{A}$ with varying $A$. As it turns out, $c_{A}$-concavity implies concavity in the usual sense, so that dual potentials can be chosen to be concave. Furthermore, when the supports of $\mu$ and $\nu$ are contained in $B_{\mathcal{X}}(0,r)$ and $B_{\mathcal{Y}}(0,r)$, respectively, one can choose $\mathcal{F}$ to be consisting of concave, uniformly bounded, and uniformly Lipschitz functions, where the uniform upper bounds on the functions themselves and Lipschitz constants scale as $r^{4}$. Now, it is not difficult to see that a version of Dudley’s entropy integral bound yields that the expectation of the first term in (10) scales as $r^{4}n^{-2/d_{x}}$. A suitable adaptation of the LCA principle (i.e., Lemma 2.1 in [HSM24]) implies that the complexity of $\mathcal{G}$ is essentially no greater than that of $\mathcal{F}$, so that the expectation of the second term in (10) also scales as $r^{4}m^{-2/d_{x}}$ (rather than $r^{4}m^{-2/d_{y}}$).
One approach to extending (7) to the unbounded support case is to mimic the approach of [MNW24], which establishes sharp rates for OT cost estimation in unbounded settings, and to derive quantitative local regularity estimates of dual potentials for a collection of costs $c_{A}$ with varying $A$. However, this approach would require imposing delicate tail conditions on $\mu,\nu$. Indeed, in OT cost estimation, [MNW24] assume that $\mu,\nu$ have sub-Weibull tails (which is stronger than $\mu,\nu$ having finite moments of all orders) and further satisfy anticoncentration properties. Furthermore, extending the LCA principle from [HSM24] to the unbounded support case appears to be highly nontrivial.
3. Main results
3.1. Upper bounds for marginals with unbounded supports
We now present our first main result that provides sample complexity upper bounds for GW estimation only under finite moment conditions. Recall the notation $\varphi_{n,m}$ from (8).
Theorem 3.1 (Upper bounds under finite moment conditions).
For given $q\in(2,\infty)$ and $M\geq 1$, we have
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})\\ m_{4q}(\mu)\vee m_{4q}(\nu)\leq M\end{subarray}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim\varphi_{n,m}+(n\wedge m)^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}\\ +(n\wedge m)^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2(d_{x}\vee d_{y})}{q}}\wedge(n\wedge m)^{\frac{d_{x}d_{y}}{2q}}\Big{\}}\log(n\wedge m),$ (11)
where the hidden constant depends only on $q,d_{x},d_{y}$ and $M$.
The theorem implies that for the $d_{x}\wedge d_{y}\geq 4$ case, when
$q>\frac{d_{x}\wedge d_{y}+2d_{x}d_{y}}{d_{x}\wedge d_{y}-2},$ (12)
the empirical estimator $D(\hat{\mu}_{n},\hat{\nu}_{m})$ achieves the rate $(n\wedge m)^{-\frac{2}{d_{x}\wedge d_{y}}}(\log(n\vee m))^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}$ that was known to hold only for compactly supported marginals. The minimal moment condition (12) scales linearly in $d_{x}\vee d_{y}$.
For the $d_{x}\wedge d_{y}<4$ case, our bound reduces to $(n\wedge m)^{-1/2}\sqrt{\log(n\wedge m)}$ when $q$ is
$q\geq d_{x}d_{y}+2.$
The rate involves the extra logarithmic factor $\sqrt{\log(n\wedge m)}$ compared with the compact support case. The extra factor is likely to be an artifact of our proof technique and caused by discretizing the domain for $A$ appearing in the variational representation from Lemma 2.2. At this moment, we are unsure whether $\sqrt{\log(n\wedge m)}$ can be removed from the bound.
Our proof is partially inspired by the recent work by [SH25], but extending directly their Theorem 1.1 to deal with (9) seems not straightforward. The proof first finds separate upper and lower bounds for $S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)$ as
$\inf_{A}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\}\leq S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\leq T_{c_{A^{\star}}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A^{\star}}}(\mu,\nu),$
where $A^{\star}$ is any matrix that achieves the infimum in the variational representation (6) and $\inf_{A}$ on the left-hand side can be reduced to the infimum over a compact subset of $\mathbb{R}^{d_{x}\times d_{y}}$ up to a sufficiently small error. Since $A^{\star}$ is a single matrix, the upper bound can be dealt with by applying Theorem 1.1 in [SH25], or its suitable modification to accommodate extra logarithmic factors (see Theorem A.1 below). For the lower bound, OT duality tells us that
$T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\geq\int f_{A}\,d(\hat{\mu}_{n}-\mu)+\int g_{A}\,d(\hat{\nu}_{m}-\nu),$
where $(f_{A},g_{A})$ are dual potentials for $(\mu,\nu)$ w.r.t. cost $c_{A}$ for each $A$. Since regularity of the mappings $A\mapsto f_{A}$ and $A\mapsto g_{A}$ is unclear in the unbounded setting, we discretize the set of $A$ and apply the maximal inequality for a finite function class from Lemma 5 in [CCK15]. Controlling the discretization error on the primal level gives the result.
The preceding theorem requires both marginals to have finite $(4q)$-th moments with $q\in(2,\infty)$. While finite $8$-th moments are (almost) necessary to ensure that, e.g., $\mathbb{E}[|S_{1}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{1}(\mu,\nu)|]\lesssim(n\wedge m)^{-1/2}$, the moment condition in Theorem 3.1 is by no means the most general. As a particular instance, we shall consider the case where one of the marginals, say $\nu$, is compactly supported. For simplicity, we assume $\nu\in\mathcal{P}(B_{\mathcal{Y}}(0,1))$ in the next proposition. Let $\bar{\varphi}_{n,m,q}$ denote the right hand side on (11).
Proposition 3.1 (Upper bounds when one marginal has compact support).
For given $q\in(2,\infty)$ and $M\geq 1$, we have
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{8\vee(2q)}(\mathcal{X})\times\mathcal{P}(B_{\mathcal{Y}}(0,1))\\ m_{8\vee(2q)}(\mu)\leq M\end{subarray}}\mathbb{E}\left[\big{|}D(\hat{\mu}_{n},\hat{\nu}_{m})-D(\mu,\nu)\big{|}\right]\lesssim\bar{\varphi}_{n,m,q},$
where the hidden constant depends only on $q,d_{x},d_{y}$ and $M$.
The proposition requires $\mu$ to have a finite $\big{(}8\vee(2q)\big{)}$-th moment instead of $4q$, while implying that the empirical estimator retains the rate $n^{-\frac{2}{d_{x}\vee d_{y}}}(\log(n\vee m))^{\mathbbm{1}_{\{d_{x}\wedge d_{y}=4\}}}$ when $d_{x}\wedge d_{y}\geq 4$ provided that $q$ satisfies (12). The reason behind this is that, in Theorem 3.1 the cost $c_{A}$ is bounded as $|c_{A}(x,y)|\lesssim 1+\|x\|^{4}+\|y\|^{4}$, while if $\nu$ is supported in $B_{\mathcal{Y}}(0,1)$, one can use a bound $|c_{A}(x,y)|\lesssim 1+\|x\|^{2}$. This implies that dual potentials for $c_{A}$ have bounded $q$-th moments if $\mu$ has a finite $(2q)$-th moment. The proof of the proposition is a reasonably minor modification of that for Theorem 3.1. In principle, one can also consider more general situations where two marginals have finite moments of different orders, which, however, will not be pursued here.
Remark 3.1 (Penalized Wasserstein alignment and Wasserstein Procrustes).
Our proof technique can be easily adapted to penalized Wasserstein alignment considered in [PSW25]. Let $\{c_{\theta}:\theta\in\Theta\}$ be a family of continuous cost functions from $\mathcal{X}\times\mathcal{Y}$ into $\mathbb{R}$ and let $\mathrm{pen}:\Theta\to\mathbb{R}$ be a (bounded) penalty function. The penalized Wasserstein alignment problem considered in [PSW25] reads as
$\inf_{\theta\in\Theta}\big{\{}T_{c_{\theta}}(\mu,\nu)+\mathrm{pen}(\theta)\big{\}},$
which encompasses the GW problem (via the variational representation) and Wasserstein Procrustes [XWLL15, GJB19]. For simplicity, we shall focus on Wasserstein Procrustes. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the (squared) Procrustes-Wasserstein distance is defined by
$\tilde{W}_{2}^{2}(\mu,\nu):=\inf_{O\in\mathcal{O}(d)}W_{2}^{2}(O_{\#}\mu,\nu)=\inf_{O\in\mathcal{O}(d)}T_{\tilde{c}_{O}}(\mu,\nu),$
where $\mathcal{O}(d)$ denotes the set of $d\times d$ orthogonal matrices and $\tilde{c}_{O}(x,y)=\|Ox-y\|^{2}=\|x\|^{2}+\|y\|^{2}-2x^{\top}Oy$.
Proposition 3.2.
Consider the above setting. Suppose that $\mu,\nu\in\mathcal{P}_{2q}(\mathbb{R}^{d})$ for some $q\in(2,\infty)$. Then,
$\mathbb{E}\big{[}|\tilde{W}_{2}^{2}(\hat{\mu}_{n},\hat{\nu}_{m})-\tilde{W}_{2}^{2}(\mu,\nu)|\big{]}\lesssim(n\wedge m)^{-\frac{2}{d}}(\log(n\vee m))^{\mathbbm{1}_{\{d=4\}}}+(n\wedge m)^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}\\ +(n\wedge m)^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2d}{q}}\wedge(n\wedge m)^{\frac{d^{2}}{2q}}\Big{\}}\log(n\wedge m).$
The hidden constant depends only on $q,d,m_{2q}(\mu)$ and $m_{2q}(\nu)$.
The proof is almost identical to Theorem 3.1, given that a version of Lemma 4.2 below (with $r^{4}$ replaced by $r^{2}$) holds for $T_{\tilde{c}_{O}}$ (see [SH25]) and hence omitted for brevity. Furthermore, a minor modification of the proof of Theorem 3.2 below yields that $(n\wedge m)^{-\frac{2}{d}}$ matches a minimax lower bound (up to a logarithmic factor) over the class of distributions supported in the unit ball.
3.2. Minimax lower bounds
For minimax lower bounds, it suffices to consider a smaller class of distributions than that in Theorem 3.1.
Theorem 3.2 (Minimax lower bounds).
The following minimax lower bound holds:
$\begin{split}&\inf_{\hat{D}_{n,m}}\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{X}}(0,1))}\mathbb{E}\left[\big{|}\hat{D}_{n,m}-D(\mu,\nu)\big{|}\right]\\ &\qquad\gtrsim(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}(\log(n\wedge m))^{-\frac{2}{d_{x}\wedge d_{y}}\mathbbm{1}_{\{d_{x}\wedge d_{y}>4\}}}\end{split}$ (13)
where the infimum is taken over all estimators of $D(\mu,\nu)$ constructed from i.i.d. samples from $\mu$ and $\nu$ of sizes $n$ and $m$, respectively, that are independent of each other. The hidden constant depends only on $d_{x}$ and $d_{y}$.
The proof builds on adapting techniques developed in [NWR22] (see also [MNW24]) and compares $\mathsf{GW}_{2,2}$ with the total vatiation and the product of the total variaton and $\chi^{2}$-divergence for random mixtures of point masses; see (21). The upper bound is straightforward from comparing $\mathsf{GW}_{2,2}$ with $W_{2}$ using Lemma 2.1. Deriving the lower comparison inequality requires more effort. First, to apply the second result in Lemma 2.1 (ii), one has to appropriately choose support points so that they are well-separated and at the same time the smallest eigenvalue of the covariance matrix of the empirical distribution is bounded away from zero; see Lemma 4.4. Second, extra care is needed to bound $\mathsf{GW}_{2,2}$ from below by the total variation because of invariance of $\mathsf{GW}_{2,2}$ under isometries; see the argument above (21).
4. Proofs
4.1. Proof of Theorem 3.1
We will use the following observation without further mentioning: $m_{p}(\mu)\leq 1+m_{p^{\prime}}(\mu)$ for any $p<p^{\prime}$.
We first prove the following auxiliary lemma.
Lemma 4.1.
Let $q\in(2,\infty)$ and $R,M\geq 1$ be given. Then,
$\sup_{\begin{subarray}{c}(\mu,\nu):~m_{4q}(\mu)\vee m_{4q}(\nu)\leq M\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right]\lesssim\varphi_{n,m},$
where the hidden constant depends only on $q,d_{x},d_{y},R$ and $M$.
The proof relies on a version of Theorem 1.1 in [SH25] stated as Theorem A.1 below. To apply the preceding theorem, we need the following result for compactly supported marginals.
Lemma 4.2.
Let $R\geq 1$ be given. Then, there exists a constant $\kappa>0$ that depends only on $d_{x},d_{y}$ and $R$ for which the following holds:
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,r))\times\mathcal{P}(B_{\mathcal{Y}}(0,r))\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right]\\ \leq\kappa r^{4}\varphi_{n,m}$ (14)
for all $r\geq 1$ and $n,m\geq 2$.
Proof of Lemma 4.2.
The proof is essentially contained in the proof of Theorem 4.2 in [ZGMS24], so we only provide an outline. For any random vector $(X,Y)$ supported in $B_{\mathcal{X}}(0,r)\times B_{\mathcal{Y}}(0,r)$,
$\mathbb{E}[c_{A}(X,Y)]=r^{4}\mathbb{E}[c_{A/r^{2}}(X/r,Y/r)].$
Since $(X/r,Y/r)$ is supported in $B_{\mathcal{X}}(0,1)\times B_{\mathcal{Y}}(0,1)$ and $\|A/r^{2}\|_{\operatorname{op}}\leq\|A\|_{\operatorname{op}}$, the left-hand side on (14) is bounded above by
$r^{4}\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right].$
Thus, it suffices to prove (14) for $r=1$.
We may assume without loss of generality that $d_{x}\leq d_{y}$, so that $d_{x}\wedge d_{y}=d_{x}$. For a sufficiently large constant $K$ that depends only on $d_{x},d_{y}$ and $R$, consider the function class
$\mathcal{F}=\Big{\{}f:B_{\mathcal{X}}(0,1)\to\mathbb{R}:\text{$f$ is concave and $\|f\|_{\infty}\vee\|f\|_{\mathrm{Lip}}\leq K$}\Big{\}},$
where $\|\cdot\|_{\infty}$ and $\|\cdot\|_{\mathrm{Lip}}$ denote the sup-norm and Lipschitz constant, respectively, i.e.,
$\|f\|_{\infty}:=\sup_{x\in B_{\mathcal{X}}(0,1)}|f(x)|\quad\text{and}\quad\|f\|_{\mathrm{Lip}}:=\sup_{\begin{subarray}{c}x,x^{\prime}\in B_{\mathcal{X}}(0,1)\\ x\neq x^{\prime}\end{subarray}}\frac{|f(x)-f(x^{\prime})|}{\|x-x^{\prime}\|}.$
Furthermore, let $\mathcal{F}^{c_{A}}$ be the set of $c_{A}$-transforms of functions in $\mathcal{F}$, i.e.,
$\mathcal{F}^{c_{A}}:=\Big{\{}f^{c_{A}}:B_{\mathcal{Y}}(0,1)\to\mathbb{R}:f\in\mathcal{F}\Big{\}}\quad\text{with}\quad f^{c_{A}}(y):=\inf_{x\in B_{\mathcal{X}}(0,1)}\{c_{A}(x,y)-f(x)\}.$
By Lemma 5.4 in [ZGMS24], versions of dual potentials for $(\mu,\nu)$ and $(\hat{\mu}_{n},\hat{\nu}_{m})$ both belong to $\mathcal{F}\times\mathcal{F}^{c_{A}}$. By duality,
$\begin{split}\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}&\leq\sup_{f\in\mathcal{F}}\left|\int f\,d(\hat{\mu}_{n}-\mu)\right|+\sup_{g\in\mathcal{F}^{c_{A}}}\left|\int g\,d(\hat{\nu}_{m}-\nu)\right|\\ &=:I+II.\end{split}$
From $\|\cdot\|_{\infty}$-entropy number estimates for $\mathcal{F}$ (cf. Corollary 2.7.10 in [vdVW96]), combined with a version of Dudley’s entropy integral bound (cf. Lemma 16 in [vLB04]), we have
$\mathbb{E}[I]\lesssim\inf_{\alpha>0}\left[\alpha+\frac{1}{\sqrt{n}}\int_{\alpha}^{1}\tau^{-d_{x}/4}\,d\tau\right]\lesssim n^{-\frac{2}{d_{x}\vee 4}}(\log n)^{\mathbbm{1}_{\{d_{x}=4\}}}$
up to a constant that depends only on $d_{x},d_{y}$ and $R$. For the second term $II$, by Lemma 2.1 in [HSM24] (or by the definition of the $c_{A}$-transform), $\|\cdot\|_{\infty}$-covering numbers for $\mathcal{F}^{c_{A}}$ are no greater than those of $\mathcal{F}$, and as such, arguing as in the previous case, we have $\mathbb{E}[II]\lesssim m^{-\frac{2}{d_{x}\vee 4}}(\log m)^{\mathbbm{1}_{\{d_{x}=4\}}}$ up to a constant that depends only on $d_{x},d_{y}$ and $R$. ∎
Proof of Lemma 4.1.
The proof applies Theorem A.1 below. The said theorem assumes that the cost function is nonnegative, so instead of working with $c_{A}$, we will work with the modified cost function
$\bar{c}_{A}(x,y):=c_{A}(x,y)+4\|x\|^{4}+16R\|x\|^{2}+4\|y\|^{4}+16R\|y\|^{2}\geq 0.$ (15)
A simple computation yields that
$\begin{split}&\mathbb{E}\left[\big{|}T_{\bar{c}_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{\bar{c}_{A}}(\mu,\nu)-T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})+T_{c_{A}}(\mu,\nu)\big{|}\right]\\ &\leq\mathbb{E}\Big{[}4\big{|}m_{4}(\hat{\mu}_{n})-m_{4}(\mu)\big{|}+4\big{|}m_{4}(\hat{\nu}_{m})-m_{4}(\nu)\big{|}\\ &\qquad+16R\big{|}\mu_{2}(\hat{\mu}_{n})-m_{2}(\mu)\big{|}+16R\big{|}m_{2}(\hat{\nu}_{m})-m_{2}(\nu)\big{|}\Big{]}\\ &\leq\left\{4\sqrt{m_{8}(\mu)}+4\sqrt{m_{8}(\nu)}+16R\sqrt{m_{4}(\mu)}+16R\sqrt{m_{4}(\nu)}\right\}(n\wedge m)^{-1/2}.\end{split}$
As such, it suffices to establish the conclusion with $c_{A}$ replaced by $\bar{c}_{A}$.
To apply Theorem A.1, we need to verify Condition (22) below. To this end, we first show that there exists $\kappa>0$ depending only on $d_{x},d_{y}$ and $R$ such that for all $r\geq 1,n,m\geq 2$, and $(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,r))\times\mathcal{P}(B_{\mathcal{Y}}(0,r))$,
$\mathbb{E}\left[\big{|}T_{\bar{c}_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{\bar{c}_{A}}(\mu,\nu)\big{|}\right]\leq\kappa r^{4}\varphi_{n,m}.$ (16)
Observe that for $(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,r))\times\mathcal{P}(B_{\mathcal{Y}}(0,r))$,
$\begin{split}&\mathbb{E}\left[\big{|}T_{\bar{c}_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{\bar{c}_{A}}(\mu,\nu)-T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})+T_{c_{A}}(\mu,\nu)\big{|}\right]\\ &\quad\leq(8r^{4}+32Rr^{2})(n\wedge m)^{-1/2}\\ &\quad\leq(8+32R)r^{4}(n\wedge m)^{-1/2}.\end{split}$
Combining Lemma 4.2, we obtain (16) with a suitable $\kappa$.
Now, observe that
$\begin{split}\bar{c}_{A}(x,y)&\leq(8\|x\|^{4}+32R\|x\|^{2})+(8\|y\|^{4}+32R\|y\|^{2})\\ &\leq\underbrace{\big{\{}(8+32R)\|x\|^{4}+32R\big{\}}}_{=:\bar{c}_{\mathcal{X}}(x)}+\underbrace{\big{\{}(8+32R)\|y\|^{4}+32R\big{\}}}_{=:\bar{c}_{\mathcal{Y}}(y)},\end{split}$
and $\|\bar{c}_{\mathcal{X}}\|_{L^{q}(\mu)}^{q}\vee\|\bar{c}_{\mathcal{Y}}\|_{L^{q}(\nu)}^{q}\lesssim 1$ up to a constant that depends only on $q,R$ and $M$. Since
$\bar{c}_{\mathcal{X}}^{-1}([0,r])=\begin{cases}B_{\mathcal{X}}\left(0,\left(\frac{r-32R}{8+32R}\right)^{1/4}\right)&r\geq 32R,\\ \varnothing&r<32R,\end{cases}$
Condition (22) holds with
$\alpha=\frac{2}{(d_{x}\wedge d_{y})\vee 4}\quad\text{and}\quad\delta=1.$
The desired result then follows from Theorem A.1. ∎
The following lemma concerning moment estimates of dual potentials will also be used.
Lemma 4.3.
Let $q,R\geq 1$ be given. Set $\bar{c}_{\mathcal{X}}(x):=(8+32R)\|x\|^{4}+32R$ and $\bar{c}_{\mathcal{Y}}(y):=(8+32R)\|y\|^{4}+32R$. For every $(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})$ and $A\in\mathbb{R}^{d_{x}\times d_{y}}$ with $\|A\|_{\operatorname{op}}\leq R$, one can find dual potentials $(f_{A},g_{A})$ for $(\mu,\nu)$ w.r.t. cost $c_{A}$ such that
$\|f_{A}\|_{L^{q}(\mu)}^{q}+\|g_{A}\|_{L^{q}(\nu)}^{q}\leq 2^{4q+3}(\|\bar{c}_{\mathcal{X}}\|_{L^{q}(\mu)}^{q}+\|\bar{c}_{\mathcal{Y}}\|_{L^{q}(\nu)}^{q}).$
Proof.
Let $\bar{c}_{A}\geq 0$ be as defined in (15). By Lemma 5.4 in [SH25], one can find dual potentials $(\bar{f}_{A},\bar{g}_{A})$ for $\bar{c}_{A}$ such that
$\|\bar{f}_{A}\|_{L^{q}(\mu)}^{q}+\|\bar{g}_{A}\|_{L^{q}(\nu)}^{q}\leq 2^{3q+3}(\|\bar{c}_{\mathcal{X}}\|_{L^{q}(\mu)}^{q}+\|\bar{c}_{\mathcal{Y}}\|_{L^{q}(\nu)}^{q}).$
We now observe that
$f_{A}(x):=\bar{f}_{A}(x)+(4\|x\|^{4}+16R\|x\|^{2})\quad\text{and}\quad g_{A}(y):=\bar{g}_{A}(y)+(4\|y\|^{4}+16R\|y\|^{2})$
are dual potentials for $c_{A}$ satisfying
$\begin{split}&\|f_{A}\|_{L^{q}(\mu)}^{q}+\|g_{A}\|_{L^{q}(\nu)}^{q}\\ &\leq 2^{q-1}\|\bar{f}_{A}\|_{L^{q}(\mu)}^{q}+2^{q-1}\|\bar{c}_{\mathcal{X}}/2\|_{L^{q}(\mu)}^{q}+2^{q-1}\|\bar{g}_{A}\|_{L^{q}(\nu)}^{q}+2^{q-1}\|\bar{c}_{\mathcal{Y}}/2\|_{L^{q}(\nu)}^{q}\\ &\leq 2^{4q+3}(\|\bar{c}_{\mathcal{X}}\|_{L^{q}(\mu)}^{q}+\|\bar{c}_{\mathcal{Y}}\|_{L^{q}(\nu)}^{q}),\end{split}$
completing the proof. ∎
We are now in position to prove Theorem 3.1.
Proof of Theorem 3.1.
Pick any $(\mu,\nu)\in\mathcal{P}_{4q}(\mathcal{X})\times\mathcal{P}_{4q}(\mathcal{Y})$ with $m_{4q}(\mu)\vee m_{4q}(\nu)\leq M$. In this proof, the notation $\lesssim$ means that an inequality holds up to a constant that depends only on $q,d_{x},d_{y}$ and $M$. Assume without loss of generality that $\mu$ and $\nu$ have mean zero. Let $\tilde{\mu}_{n}$ and $\tilde{\nu}_{m}$ be the centered versions of $\hat{\mu}_{n},\hat{\nu}_{m}$, respectively. We first observe that
$D(\hat{\mu}_{n},\hat{\nu}_{m})=D(\tilde{\mu}_{n},\tilde{\nu}_{m})=S_{1}(\tilde{\mu}_{n},\tilde{\nu}_{m})+S_{2}(\tilde{\mu}_{n},\tilde{\nu}_{m}).$
By direct calculations, it is not difficult to see that
$\begin{split}\mathbb{E}\big{[}|S_{1}(\tilde{\mu}_{n},\tilde{\nu}_{m})-S_{1}(\mu,\nu)|\big{]}&\lesssim(n\wedge m)^{-1/2},\\ \mathbb{E}\left[|S_{2}(\tilde{\mu}_{n},\tilde{\nu}_{m})-S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})\big{|}\right]&\lesssim(n\wedge m)^{-1/2}.\end{split}$ (17)
These estimates follow from tedious but straightforward computations, having in mind, e.g., $\mathbb{E}[\|\bar{X}\|^{4}]\lesssim n^{-2}$ for $\bar{X}=n^{-1}\sum_{i=1}^{n}X_{i}$.
Let $A^{\star}$ be a matrix with $\|A^{\star}\|_{\operatorname{op}}\leq\sqrt{m_{2}(\mu)m_{2}(\nu)}/2$ that achieves the infimum in the variational representation (6),
$S_{2}(\mu,\nu)=32\|A^{\star}\|_{\mathrm{F}}^{2}+T_{c_{A^{\star}}}(\mu,\nu).$
Applying the variational representation to $S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})$, one has
$S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\leq T_{c_{A^{\star}}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A^{\star}}}(\mu,\nu).$
To find a lower bound, define the event
$\mathcal{E}_{n,m}=\big{\{}m_{2}(\hat{\mu}_{n})\leq m_{2}(\mu)+1\big{\}}\cap\{m_{2}(\hat{\nu}_{m})\leq m_{2}(\nu)+1\big{\}}.$
By Markov’s inequality, we have
$\begin{split}\mathbb{P}(\mathcal{E}_{n,m}^{c})&\leq\mathbb{P}\big{(}m_{2}(\hat{\mu}_{n})-m_{2}(\mu)\geq 1\big{)}+\mathbb{P}\big{(}m_{2}(\hat{\nu}_{m})-m_{2}(\nu)\geq 1\big{)}\\ &\leq\frac{m_{4}(\mu)}{n}+\frac{m_{4}(\nu)}{m}\\ &\leq\big{(}m_{4}(\mu)+m_{4}(\nu)\big{)}(n\wedge m)^{-1}.\end{split}$
Observe that on the event $\mathcal{E}_{n}$, by Lemma 2.2, the following variational representation holds:
$S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})=\inf_{A\in\mathcal{A}}\big{\{}32\|A\|_{\mathrm{F}}^{2}+T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})\big{\}},$
where $\mathcal{A}:=\big{\{}A\in\mathbb{R}^{d_{x}\times d_{y}}:\|A\|_{\operatorname{op}}\leq R\big{\}}$ with $R:=\sqrt{(m_{2}(\mu)+1)(m_{2}(\nu)+1)}/2$. An analogous variational representation holds for $S_{2}(\mu,\nu)$. As such, on the event $\mathcal{E}_{n}$,
$S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\geq\inf_{A\in\mathcal{A}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}.$
We further discretize the set $\mathcal{A}$. Observe that
$\begin{split}|c_{A}(x,y)-c_{B}(x,y)|&\leq 32|x^{\top}(A-B)y|\\ &\leq 32\|x\|\|y\|\|A-B\|_{\operatorname{op}}\\ &\leq 16(\|x\|^{2}+\|y\|^{2})\|A-B\|_{\operatorname{op}}.\end{split}$
For $\epsilon>0$, let $\mathcal{N}_{\epsilon}$ be an $\epsilon$-net for $\mathcal{A}$ w.r.t. $\|\cdot\|_{\operatorname{op}}$, so that for any $A\in\mathcal{A}$, there exists $B\in\mathcal{N}_{\epsilon}$ such that
$\|A-B\|_{\operatorname{op}}\leq\epsilon.$
By volumetric argument, it is not difficult to see that
$|\mathcal{N}_{\epsilon}|\leq\left(\frac{3R}{\epsilon}\right)^{d_{x}d_{y}}$
for all $0<\epsilon<1$. We shall choose
$\epsilon=\epsilon_{n,m}:=(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}.$
By construction,
$\begin{split}\inf_{A\in\mathcal{A}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}&\geq\min_{A\in\mathcal{N}_{\epsilon_{n,m}}}\big{\{}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{\}}\\ &\quad-16\big{(}m_{2}(\hat{\mu}_{n})+m_{2}(\hat{\nu}_{m})+m_{2}(\mu)+m_{2}(\nu)\big{)}\epsilon_{n,m}.\end{split}$
For each $A\in\mathcal{A}$, let $(f_{A},g_{A})$ be dual potentials for $(\mu,\nu)$ w.r.t. cost $c_{A}$. By duality,
$T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\geq\int f_{A}\,d(\hat{\mu}_{n}-\mu)+\int g_{A}\,d(\hat{\nu}_{m}-\nu).$
Outside $\mathcal{E}_{n}$, we use the following crude bound:
$\begin{split}S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)&\geq-4m_{4}(\hat{\mu}_{n})-16Rm_{2}(\hat{\mu}_{n})-4m_{4}(\hat{\nu}_{m})-16Rm_{2}(\hat{\nu}_{m})\\ &\quad-4m_{4}(\mu)-16Rm_{2}(\mu)-4m_{4}(\nu)-16Rm_{2}(\nu).\end{split}$
Now, we observe that
$\begin{split}\mathbb{E}\left[\big{|}S_{2}(\hat{\mu}_{n},\hat{\nu}_{m})-S_{2}(\mu,\nu)\big{|}\right]&\lesssim\mathbb{E}\left[\big{|}T_{c_{A^{\star}}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A^{\star}}}(\mu,\nu)\big{|}\right]\\ &\qquad+\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int f_{A}\,d(\hat{\mu}_{n}-\mu)\right|\right]\\ &\qquad+\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int g_{A}\,d(\hat{\nu}_{m}-\nu)\right|\right]\\ &\qquad+\mathbb{E}\big{[}\big{(}m_{4}(\hat{\mu}_{n})+m_{4}(\hat{\nu}_{m})\big{)}\mathbbm{1}_{\mathcal{E}_{n,m}^{c}}\big{]}+(n\wedge m)^{-\frac{2}{(d_{x}\wedge d_{y})\vee 4}}.\end{split}$
By the Cauchy-Schwarz inequality, the fourth term on the right-hand side is $\lesssim(n\wedge m)^{-1/2}$, and Lemma 4.1 yields that the first term is $\lesssim\varphi_{n,m}$.
It remains to find upper bounds on
$\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int f_{A}\,d(\hat{\mu}_{n}-\mu)\right|\right]\quad\text{and}\quad\mathbb{E}\left[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}\left|\int g_{A}\,d(\hat{\nu}_{m}-\mu)\right|\right].$
Set
$Y_{A}=\left|\int f_{A}\,d(\hat{\mu}_{n}-\mu)\right|\quad\text{and}\quad W_{A}=\left|\int g_{A}\,d(\hat{\nu}_{m}-\nu)\right|.$
Having in mind the fact that $\mathcal{N}_{\epsilon_{n,m}}$ is a finite set, we apply Lemma 8 in [CCK15] (restated in Lemma A.1 below) to find bounds on $\mathbb{E}[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}|Y_{A}|]$ and $\mathbb{E}[\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}|W_{A}|]$. By Lemma 4.3, one can choose versions of $f_{A}$ and $g_{A}$ in such a way that
$\sup_{A\in\mathcal{A}}\Big{(}\|f_{A}\|_{L^{q}(\mu)}^{q}\vee\|g_{A}\|_{L^{q}(\nu)}^{q}\Big{)}\lesssim 1.$
This implies $\sup_{A\in\mathcal{A}}\|f_{A}\|_{L^{2}(\mu)}^{2}\lesssim 1$ and
$\mathbb{E}\left[\max_{1\leq i\leq n}\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}f_{A}^{2}(X_{i})\right]\lesssim(n|\mathcal{N}_{\epsilon_{n,m}}|)^{2/q}$
Recalling that $|\mathcal{N}_{\epsilon_{n,m}}|\lesssim\epsilon_{n,m}^{-d_{x}d_{y}}\lesssim(n\wedge m)^{2(d_{x}\vee d_{y})}\wedge(n\wedge m)^{\frac{d_{x}d_{y}}{2}}$, by Lemma A.1, we have
$\mathbb{E}\Big{[}\max_{A\in\mathcal{N}_{\epsilon_{n,m}}}|Y_{A}|\Big{]}\lesssim n^{-\frac{1}{2}}\sqrt{\log(n\wedge m)}+n^{\frac{1-q}{q}}\Big{\{}(n\wedge m)^{\frac{2(d_{x}\vee d_{y})}{q}}\wedge(n\wedge m)^{\frac{d_{x}d_{y}}{2q}}\Big{\}}\log(n\wedge m).$
A similar estimate holds for $W_{A}$. Putting everything together, we obtain the desired conclusion. This completes the proof of the theorem. ∎
4.2. Proof of Proposition 3.1
The proof is a modification to that of Theorem 3.1. First, we observe that, when $(X,Y)$ is supported in $B_{\mathcal{X}}(0,r)\times B_{\mathcal{Y}}(0,1)$,
$\mathbb{E}[c_{A}(X,Y)]=r^{2}\mathbb{E}[c_{A/r}(X/r,Y)],$
and $(X/r,Y)$ is supported in $B_{\mathcal{X}}(0,1)\times B_{\mathcal{Y}}(0,1)$. As such, one has
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,r))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right]\\ \leq\kappa r^{2}\varphi_{n,m}.$
The modified cost is now replaced with
$\bar{c}_{A}(x,y)=c_{A}(x,y)+(8+16R)\|x\|^{2}+16R$
which is nonnegative on $\mathcal{X}\times B_{\mathcal{Y}}(0,1)$ and uppper bounded by
$\bar{c}_{A}(x,y)\leq\underbrace{(16+32R)\|x\|^{2}}_{=:\bar{c}_{\mathcal{X}}(x)}+\underbrace{32R}_{=:\bar{c}_{\mathcal{Y}}(y)}.$
Applying Theorem A.1 with $\mathcal{Y}=B_{\mathcal{Y}}(0,1)$ and $\mathcal{B}_{\mathcal{Y}}(r)=B_{\mathcal{Y}}(0,1)$ for all $r\geq 1$, we have, for given $q\in(2,\infty)$ and $R,M\geq 1$,
$\sup_{\begin{subarray}{c}(\mu,\nu)\in\mathcal{P}_{2q}(\mathcal{X})\times\mathcal{P}(B_{\mathcal{Y}}(0,1))\\ \|A\|_{\operatorname{op}}\leq R\end{subarray}}\mathbb{E}\left[\big{|}T_{c_{A}}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c_{A}}(\mu,\nu)\big{|}\right]\\ \lesssim\varphi_{n,m},$
up to a constant that depends only on $q,d_{x},d_{y},R$ and $M$.
Having in mind that the inequalities in (17) hold under finite $8$-th moments, the rest of the proof is completely analogous to Theorem 3.1. ∎
4.3. Proof of Theorem 3.2
We first define some notations. Let $P,Q$ be probability measures defined on a common measurable space.
  • (Total variation)
    $d_{\mathrm{TV}}(P,Q):=\sup_{A}|P(A)-Q(A)|;$
  • ($\chi^{2}$-divergence)
    $\chi^{2}(P,Q):=\begin{cases}\int\left(\frac{dP}{dQ}-1\right)^{2}\,dQ&\text{if $P\ll Q$},\\ \infty&\text{otherwise}.\end{cases}$
We will use the following properties of the total variation and $\chi^{2}$-divergence:
$d_{\mathrm{TV}}(P,Q)$ $=\inf_{\pi\in\Pi(P,Q)}\mathbb{P}_{(X,Y)\sim\pi}(X\neq Y),$ (18)
$d_{\mathrm{TV}}(P,Q)$ $\leq\sqrt{\chi^{2}(P,Q)},$ (19)
$\chi^{2}(P^{n},Q^{n})$ $=\big{(}1+\chi^{2}(P,Q)\big{)}^{n}-1,$ (20)
where $P^{n}=\otimes_{i=1}^{n}P$ and $Q^{n}=\otimes_{i=1}^{n}Q$. The last two properties follow directly from the definitions; see [Tsy09, p. 86 and p. 90]. The first property is also well-known.
We first prove two auxiliary lemmas that will be used in the proof of Theorem 3.2.
Lemma 4.4.
There exists a constant $c>0$ depending on $d$ only such that, for every sufficiently large positive integer $k$, there exists a set $\{x_{1},\ldots,x_{k}\}\subset B_{\mathbb{R}^{d}}(0,1)$ that satisfies $\left\lVert x_{i}-x_{j}\right\rVert\geq ck^{-1/d}$ for all $i\neq j$, and the covariance matrix $\Sigma_{\mu}$ of the distribution $\mu=k^{-1}\sum_{i=1}^{k}\delta_{x_{i}}$ satisfies $\lambda_{\min}(\Sigma_{\mu})\geq c$.
Proof.
We start with verifying that one can choose points $\{x_{1},\ldots,x_{k}\}$ such that $\mu$ has mean zero. For a given integer $k$, consider a maximal set of points $\{x_{1}^{\prime},\ldots,x_{1}^{\prime}\}$ inside $B_{\mathbb{R}^{d}}(0,1/3)$ such that $\|x_{i}^{\prime}-x_{j}^{\prime}\|>\gamma_{d}k^{-1/d}$ for all $i\neq j$, where $\gamma_{d}$ is a small positive constant that depends only on $d$. Let $x_{i}:=x_{i}^{\prime}-\bar{x}^{\prime}$ with $\bar{x}^{\prime}=k^{-1}\sum_{i=1}^{k}x_{i}^{\prime}$, then $\{x_{1},\ldots,x_{k}\}\subset B_{\mathbb{R}^{d}}(0,2/3)$, $\left\lVert x_{i}-x_{j}\right\rVert>\gamma_{d}k^{-1/d}$ for all $i\neq j$, and the distribution $\mu=k^{-1}\sum_{i=1}^{k}\delta_{x_{i}}$ has mean zero.
Observe that
$\lambda_{\min}(\Sigma_{\mu})=\min_{\|v\|=1}v^{\top}\Sigma_{\mu}v=\min_{\|v\|=1}\frac{1}{k}\sum_{i=1}^{k}(v^{\top}x_{i})^{2}\,.$
Let $v$ be an arbitrary unit vector in $\mathbb{R}^{d}$. Consider the set of disjoint balls of radius $r_{k}:=\frac{\gamma_{d}}{2}k^{-1/d}$ centered at $x_{i}$. Recall that
$\mathrm{Vol}(B_{\mathbb{R}^{d}}(0,r_{k}))=\alpha_{d}r_{k}^{d}=\alpha_{d}(\gamma_{d}/2)^{d}k^{-1},$
where $\alpha_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$. Let
$S_{\delta}:=\{x\in\mathbb{R}^{d}:|v^{\top}x|\leq\delta\}$
for some $\delta>0$ to be chosen later. Let $N_{\delta}=|\{x_{1},\dots,x_{k}\}\cap S_{\delta}|$. The $N_{\delta}$ disjoint balls corresponding to these points are all contained in $S_{\delta+r_{k}}\cap B_{\mathbb{R}^{d}}(0,1)$ (for sufficiently large $k$). By comparing volumes, we have
$N_{\delta}\cdot\left(\alpha_{d}\left(\gamma_{d}\over 2\right)^{d}k^{-1}\right)\leq 2(\delta+r_{k})\alpha_{d-1},$
which implies
$N_{\delta}\leq k\cdot\frac{2(\delta+r_{k})\alpha_{d-1}}{\alpha_{d}(\gamma_{d}/2)^{d}}\leq k\cdot\frac{3\delta\alpha_{d-1}}{\alpha_{d}(\gamma_{d}/2)^{d}}$
for $k$ large enough. Let $K_{d,\gamma_{d}}:=\frac{3\alpha_{d-1}}{\alpha_{d}(\gamma_{d}/2)^{d}}$, and choose $\delta=\frac{1}{2K_{d,\gamma_{d}}}$, which yields $N_{\delta}\leq k/2$. Since at most $k/2$ points are inside $S_{\delta}$, at least $k/3$ points must be outside $S_{\delta}$. For any point $x_{i}$ outside $S_{\delta}$, we have $(v^{\top}x_{i})^{2}>\delta^{2}$. Therefore,
$\frac{1}{k}\sum_{i=1}^{k}(v^{\top}x_{i})^{2}\geq\frac{1}{k}\sum_{x_{i}\notin S_{\delta}}(v^{\top}x_{i})^{2}>\frac{k/3}{k}\delta^{2}=\frac{\delta^{2}}{3}\,.$
We conclude that
$\lambda_{\min}(\Sigma_{\mu})\geq\frac{\delta^{2}}{3}=\frac{1}{3}\left(\frac{1}{2K_{d,\gamma_{d}}}\right)^{2}=\frac{1}{12}\left(\frac{\alpha_{d}(\gamma_{d}/2)^{d}}{3\alpha_{d-1}}\right)^{2}.$
This completes the proof.
Lemma 4.5.
Suppose $\nu=\frac{1}{2}\delta_{-1}+\frac{1}{2}\delta_{1}$ and $\mu=(\frac{1}{2}+\epsilon)\delta_{-1}+(\frac{1}{2}-\epsilon)\delta_{1}$. Then for sufficiently small $\epsilon>0$, we have $D(\mu,\nu)=32\epsilon(1-\epsilon)$.
Proof.
By definition,
$D(\mu,\nu)=\inf_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{(X,Y,X^{\prime}Y^{\prime})\sim\pi\otimes\pi}\left[|X-X^{\prime}|^{4}+|Y-Y^{\prime}|^{4}-2|X-X^{\prime}|^{2}|Y-Y^{\prime}|^{2}\right].$
Pick any $\pi\in\Pi(\mu,\nu)$ and let $(X,Y)\sim\pi$ and $(X^{\prime},Y^{\prime})\sim\pi$ be independent. Observe that
$\begin{split}\mathbb{E}\left[|X-X^{\prime}|^{4}\right]&=16\mathbb{P}(X\neq X^{\prime})\\ &=16\left(1-\left(\left(\frac{1}{2}+\epsilon\right)^{2}+\left(\frac{1}{2}-\epsilon\right)^{2}\right)\right)\\ &=8-32\epsilon^{2},\\ \mathbb{E}\left[|Y-Y^{\prime}|^{4}\right]&=16\mathbb{P}(Y\neq Y^{\prime})=8,\quad\text{and}\\ \mathbb{E}\left[|X-X^{\prime}|^{2}|Y-Y^{\prime}|^{2}\right]&=\mathbb{E}\left[(2-2XX^{\prime})(2-2YY^{\prime})\right]\\ &=4\mathbb{E}\left[(1-XX^{\prime})(1-YY^{\prime})\right]\\ &=4\left(1-\mathbb{E}[X]^{2}-\mathbb{E}[Y]^{2}+\mathbb{E}[XY]^{2}\right)\\ &=4\left(1-4\epsilon^{2}+\mathbb{E}[XY]^{2}\right)\end{split}$
so that
$\begin{split}D(\mu,\nu)&=\inf_{\pi\in\Pi(\mu,\nu)}\left\{16-32\epsilon^{2}-2\cdot 4(1-4\epsilon^{2}+\mathbb{E}_{\pi}[XY]^{2})\right\}\\ &=8-8\sup_{\pi\in\Pi(\mu,\nu)}\mathbb{E}_{\pi}[XY]^{2}\,.\end{split}$
To compute the supremum, we note that any coupling $\pi$ between $\mu$ and $\nu$ is determined uniquely by the single parameter $a:=\pi(\{-1,-1\})$, where the valid range for $a$ is $\epsilon,1/2]$. Then, by direct computation,
$\begin{split}\mathbb{E}_{\pi}[XY]&=a+(-1)\left(\frac{1}{2}-a\right)+(-1)\left(\frac{1}{2}+\epsilon-a\right)+(a-\epsilon)\\ &=4a-1-2\epsilon\,.\end{split}$
We just need to minimize $(4a-1-2\epsilon)^{2}$ over $a\in[\epsilon,1/2]$. Clearly, the maximum occurs either at $a=\epsilon$ or $a=1/2$. At $a=\epsilon$, we have $(4a-1-2\epsilon)^{2}=(2\epsilon-1)^{2}$. At $a=1/2$, we also have $(4a-1-2\epsilon)^{2}=(2\epsilon-1)^{2}$. So the maximum value is $(2\epsilon-1)^{2}=1-4\epsilon+4\epsilon^{2}$. We conclude that
$\begin{split}D(\mu,\nu)&=8-8+32\epsilon-32\epsilon^{2}\\ &=32\epsilon(1-\epsilon),\end{split}$
completing the proof. ∎
We are now in position to prove Theorem 3.2.
Proof of Theorem 3.2.
Let $\mathcal{M}_{n,m}$ denote the left-hand side on (13). In this proof, the notation $\lesssim$ means that an inequality holds up to a constant that depends only on $d_{x}$ and $d_{y}$. By symmetry, we may assume without loss of generality that $d_{x}\leq d_{y}$ and $n\leq m$. We divide the proof into two steps.
Step 1. First, we shall establish that the parametric lower bound $n^{-1/2}$ always holds. Consider first the $d_{x}=d_{y}=1$ case. Let $\mu_{0}=\nu=\frac{1}{2}\delta_{-1}+\frac{1}{2}\delta_{1}$ and $\mu_{1}=(\frac{1}{2}+\epsilon)\delta_{-1}+(\frac{1}{2}-\epsilon)\delta_{1}$. Then, by (19) and (20),
$d_{\mathrm{TV}}(\mu_{0}^{n},\mu_{1}^{n})\leq\sqrt{\chi^{2}(\mu_{0}^{n},\mu_{1}^{n})}=\sqrt{(1+\chi^{2}(\mu_{0},\mu_{1}))^{n}-1}\,.$
By the definition of the $\chi^{2}$-divergence,
$\begin{split}\chi^{2}(\mu_{0},\mu_{1})&=\sum_{x\in\{-1,1\}}\frac{(\mu_{0}(x)-\mu_{1}(x))^{2}}{\mu_{0}(x)}\\ &=\frac{\epsilon^{2}}{1/2}+\frac{\epsilon^{2}}{1/2}=4\epsilon^{2}\,.\end{split}$
We set $\epsilon=cn^{-1/2}$ for some small positive constant $c$, so that $\chi^{2}(\mu_{0},\mu_{1})=4c^{2}/n$ and
$\begin{split}d_{\mathrm{TV}}(\mu_{0}^{n},\mu_{1}^{n})&\leq\sqrt{\left(1+\frac{4c^{2}}{n}\right)^{n}-1}\\ &\leq\sqrt{e^{4c^{2}}-1},\end{split}$
where we used the inequality $(1+t/n)^{n}\leq e^{t}$ for $t>0$. We conclude that $d_{\mathrm{TV}}(\mu_{0}^{n},\mu_{1}^{n})<1$ if we choose $c$ to be sufficiently small.
Recalling Lemma 4.5, set
$\theta_{0}:=D(\mu_{0},\nu)=0\quad\text{and}\quad\theta_{1}:=D(\mu_{1},\nu)=32\epsilon(1-\epsilon).$
Observe that
$d_{\mathrm{TV}}(\mu_{0}^{n}\otimes\nu^{m},\mu_{1}^{n}\otimes\nu^{m})=d_{\mathrm{TV}}(\mu_{0}^{n},\mu_{1}^{n})$
which is bounded away from $1$ as argued above. By Le Cam’s two-point argument [Wu20, Theorem 9.4],
$\begin{split}\mathcal{M}_{n,m}\geq\inf_{\hat{\theta}}\max_{i\in\{0,1\}}\mathbb{E}_{(\mu_{i},\nu)}\left[\big{|}\hat{\theta}-\theta_{i}\big{|}\right]&\geq\frac{\left\lvert\theta_{1}-\theta_{0}\right\rvert}{4}\left(1-d_{\mathrm{TV}}(\mu_{0}^{n},\mu_{1}^{n})\right)\\ &\gtrsim 32\epsilon(1-\epsilon)\\ &\gtrsim 32cn^{-1/2}(1-cn^{-1/2})\\ &\gtrsim n^{-1/2}\,.\end{split}$
For general $d_{x}$ and $d_{y}$, by considering $\mu$ and $\nu$ such that the last $d_{x}-1$ and $d_{y}-1$ coordinates of $X\sim\mu$ and $Y\sim\nu$, respectively, are degenerate to 0, one can see that $\mathcal{M}_{n,m}\gtrsim n^{-1/2}$.
Step 2. Consider the $4<d_{x}\leq d_{y}$ case. By considering $\nu$ such that the last $d_{y}-d_{x}$ coordinates of $Y\sim\nu$ are degenerate to 0, it suffices to consider the $d_{x}=d_{y}=:d$ case.
For a given integer $k$, let $\{x_{1},\ldots,x_{k}\}$ be a set constructed in Lemma 4.4. Let $F$ be a random function uniformly distributed over the collection of bijections from $k]:=\{1,\dots,k\}$ onto $\{x_{1},\ldots,x_{k}\}$. Let $\mathfrak{u}$ be the uniform distribution on $[k]$ and $\mathfrak{q}$ be any distribution on $[k]$. Since the support of $F_{\#}\mathfrak{u}$ (and $F_{\#}\mathfrak{q}$) is contained in the unit ball, Lemma 2.1 (i) combined with the first inequality in Lemma 2.1 (ii) yields
$\mathsf{GW}_{2,2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\leq 4\mathsf{GW}_{2,1}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\lesssim W_{2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\,.$
Since $d>4$, one can invoke Proposition 9 in [NWR22] to conclude that there exists a constant $C_{2}>0$ depending only on $d$ such that
$\mathsf{GW}_{2,2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\leq C_{2}k^{-1/d}(\chi^{2}(\mathfrak{q},\mathfrak{u}))^{1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u})^{\frac{1}{2}-\frac{2}{d}}$
with probability at least .9.9.
On the other hand, the second inequality in Lemma 2.1 (ii) combined with Lemma 4.4 tells us that
$\begin{split}\mathsf{GW}_{2,2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})&\gtrsim\inf_{U\in E(d)}W_{2}(F_{\#}\mathfrak{q},U_{\#}(F_{\#}\mathfrak{u}))\\ &=\inf_{U\in E(d)}W_{2}(F_{\#}\mathfrak{q},(U\circ F)_{\#}\mathfrak{u})\\ &\geq\inf_{U\in E(d)}W_{1}(F_{\#}\mathfrak{q},(U\circ F)_{\#}\mathfrak{u}),\end{split}$
where we note that $F_{\#}\mathfrak{u}=k^{-1}\sum_{i=1}^{k}\delta_{x_{i}}$. Let $U$ be any element in $E(d)$ and $\pi$ be any coupling between $F_{\#}\mathfrak{q}$ and $(U\circ F)_{\#}\mathfrak{u}$. Write $y_{i}=Ux_{i}$ for $i=1,\ldots,k$. Since $U$ is an isometry, $\left\lVert y_{i}-y_{j}\right\rVert\geq ck^{-1/d}$ for $i\neq j$. For each $x_{i}$, there is at most one $y_{j}$ that satisfies $\left\lVert x_{i}-y_{j}\right\rVert\leq\frac{1}{3}ck^{-1/d}$. Similarly, for each $y_{i}$, there is at most one $x_{j}$ that satisfies $\left\lVert x_{j}-y_{i}\right\rVert\leq\frac{1}{3}ck^{-1/d}$. Hence, there exists a bijection $f$ between $\{y_{1},\ldots,y_{k}\}$ and $\{x_{1},\ldots,x_{k}\}$ such that $f(y_{i})=x_{j}$ whenever $\left\lVert y_{i}-x_{j}\right\rVert\leq\frac{1}{3}ck^{-1/d}$. This argument yields
$\left\lVert x-y\right\rVert\geq\frac{1}{3}ck^{-1/d}\mathbbm{1}\{x\neq f(y)\}$
for every $x\in\{x_{1},\ldots,x_{k}\}$ and $y\in\{y_{1},\ldots,y_{k}\}$. Hence
$\begin{split}\int\left\lVert x-y\right\rVert\,d\pi(x,y)&\geq\frac{1}{3}ck^{-1/d}\mathbb{P}_{\pi}(X\neq f(Y))\\ &\geq\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(F_{\#}\mathfrak{q},(f\circ U\circ F)_{\#}\mathfrak{u})\\ &=\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\\ &=\frac{1}{3}ck^{-1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u}),\end{split}$
where the second inequality follows from (18) and the penultimate step follows from the fact that $f\circ U$ is a bijection from $\{x_{1},\ldots,x_{k}\}$ onto itself and that $F_{\#}\mathfrak{u}$ has the uniform distribution on $\{x_{1},\ldots,x_{k}\}$. Thus we have proved
$\mathsf{GW}_{2,2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\gtrsim k^{-1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u})$
almost surely.
In summary, there exist constants $C_{1},C_{2}>0$ depending only on $d$ such that
$C_{1}k^{-1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u})\leq D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\leq C_{2}k^{-1/d}(\chi^{2}(\mathfrak{q},\mathfrak{u}))^{1/d}d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u})^{\frac{1}{2}-\frac{2}{d}},$ (21)
where the lower bound holds almost surely and the upper bound holds with probability at least .9.9.
Set $\Delta_{d}=\frac{1}{16}C_{1}k^{-1/d}$. Following [NWR22], let $\mathcal{D}_{k}$ be the subset of probability distributions $\mathfrak{q}$ on $[k]$ satisfying $\chi^{2}(\mathfrak{q},\mathfrak{u})\leq 9$. Denote by $\mathcal{D}_{k,\delta}^{-}$ the subset of $\mathcal{D}_{k}$ satisfying $d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u})\leq\delta$ and by $\mathcal{D}_{k}^{+}$ the subset of $\mathcal{D}_{k}$ satisfying $d_{\mathrm{TV}}(\mathfrak{q},\mathfrak{u})\geq 1/4$. If $\delta\leq\left(\frac{C_{1}}{144C_{2}}\right)^{\frac{1}{1/2-2/d}}$, then for any $\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}$, it holds that $D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\leq\Delta_{d}$ with probability at least .9.9. Also, for any $\mathfrak{q}\in\mathcal{D}_{m}^{+}$, it holds that $D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\geq 3\Delta_{d}$ almost surely.
Let $\hat{D}_{n,m}$ be any estimator for $D(\cdot,\cdot)$. Given $F$, generate
$X_{1},\dots,X_{n}\sim F_{\#}\mathfrak{q}\quad\text{and}\quad Y_{1},\dots,Y_{m}\sim F_{\#}\mathfrak{u}$
independently. Denote by $\mathbb{P}_{\mathfrak{q}}$ the unconditional law of $(X_{1},\dots,X_{n},Y_{1},\dots,Y_{m})$ and by $\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}$ the conditional law given $F$. Consider the test
$\psi:=\mathbbm{1}\{\hat{D}^{1/2}_{n,m}\leq 2\Delta_{d}\}.$
Define the event $A=\{|\hat{D}_{n,m}^{1/2}-D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})|\geq\Delta_{d}\}$. We obtain, for any $\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}$,
[…]
and for $\mathfrak{q}\in\mathcal{D}_{k}^{+}$, we have
$\begin{split}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}&\geq\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}\{\hat{D}_{n,m}^{1/2}\leq 2\Delta_{d}\}\cap\{D^{1/2}(F_{\#}\mathfrak{q},F_{\#}\mathfrak{u})\geq 3\Delta_{d}\}\big{)}\big{]}\\ &=\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}\hat{D}_{n,m}^{1/2}\leq 2\Delta_{d}\big{)}\big{]}\\ &=\mathbb{P}_{\mathfrak{q}}(\psi=1).\end{split}$
Conclude that
$\begin{split}&\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))}\mathbb{P}\big{(}|\hat{D}_{n,m}^{1/2}-D^{1/2}(\mu,\nu)|\geq\Delta_{d}\big{)}\\ &\quad\geq\frac{1}{2}\left(\sup_{\mathfrak{q}\in\mathcal{D}_{k}^{+}}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}+\sup_{\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}}\mathbb{E}\big{[}\mathbb{P}_{F_{\#}\mathfrak{q},F_{\#}\mathfrak{u}}\big{(}A)\big{]}\right)\\ &\quad\geq\frac{1}{2}\left(\sup_{{\mathfrak{q}\in\mathcal{D}_{k}^{+}}}\mathbb{P}_{\mathfrak{q}}(\psi=1)+\sup_{\mathfrak{q}\in\mathcal{D}_{k,\delta}^{-}}\mathbb{P}_{\mathfrak{q}}(\psi=0)\right)-.1\,.\end{split}$
Now, choosing $k=\lceil C\delta^{-1}n\log n\rceil$ for a sufficiently large constant $C$ and applying Proposition 10 of [NWR22] yields
$\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))}\mathbb{E}\left[\big{|}\hat{D}_{n,m}^{1/2}-D^{1/2}(\mu,\nu)\big{|}\right]\geq 0.8\Delta_{d}\gtrsim(n\log n)^{-1/d},$
which, by Jensen’s inequality, implies
$\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))}\mathbb{E}\left[\left(\hat{D}_{n,m}^{1/2}-D^{1/2}(\mu,\nu)\right)^{2}\right]\gtrsim(n\log n)^{-2/d}.$
Finally, by the elementary inequality $|a^{2}-b^{2}|\geq(|a|-|b|)^{2}$, we conclude that
$\sup_{(\mu,\nu)\in\mathcal{P}(B_{\mathcal{X}}(0,1))\times\mathcal{P}(B_{\mathcal{Y}}(0,1))}\mathbb{E}\left[\big{|}\hat{D}_{n,m}-D(\mu,\nu)\big{|}\right]\gtrsim(n\log n)^{-2/d}\,.$
This completes the proof. ∎
Appendix A Technical tools
A.1. A version of Theorem 1.1 in [SH25]
The following is a version of Theorem 1.1 in [SH25] that can accommodate extra logarithmic factors at the expense of less tight moment conditions. The theorem below holds for general Polish spaces $\mathcal{X}$ and $\mathcal{Y}$.
Theorem A.1.
Let $c:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}_{+}$ be a nonnegative lower semicontinuous function such that there exist nonnegative measurable functions $c_{\mathcal{X}}:\mathcal{X}\to\mathbb{R}_{+}$ and $c_{\mathcal{Y}}:\mathcal{Y}\to\mathbb{R}_{+}$ with $c\leq c_{\mathcal{X}}\oplus c_{\mathcal{Y}}$. Set $\mathcal{B}_{\mathcal{X}}(r)=c_{\mathcal{X}}^{-1}([0,r])$ and $\mathcal{B}_{\mathcal{Y}}(r)=c_{\mathcal{Y}}^{-1}([0,r])$ for $r>0$. Suppose that there exist constants $\kappa>0,\alpha\in(0,1/2]$ and $\delta>0$ such that
$\sup_{(\mu,\nu)\in\mathcal{P}(\mathcal{B}_{\mathcal{X}}(r))\times\mathcal{P}(\mathcal{B}_{\mathcal{Y}}(r))}\mathbb{E}\left[\big{|}T_{c}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c}(\mu,\nu)\big{|}\right]\leq\kappa r(n\wedge m)^{-\alpha}\big{(}\log_{+}(n\vee m)\big{)}^{\delta}$ (22)
for all $r\geq 1$ and $n,m\in\mathbb{N}$, where $\log_{+}(k)=\log(1+k)$. For given $\epsilon>0$ and $M\geq 1$, we set
$\mathcal{Q}_{\epsilon,M}=\Big{\{}(\mu,\nu)\in\mathcal{P}(\mathcal{X})\times\mathcal{P}(\mathcal{Y}):\|c_{\mathcal{X}}\|_{L^{2+\epsilon}(\mu)}^{2+\epsilon}\vee\|c_{\mathcal{Y}}\|_{L^{2+\epsilon}(\nu)}^{2+\epsilon}\leq M\Big{\}}.$
Then,
$\sup_{(\mu,\nu)\in\mathcal{Q}_{\epsilon,M}}\mathbb{E}\left[\big{|}T_{c}(\hat{\mu}_{n},\hat{\nu}_{m})-T_{c}(\mu,\nu)\big{|}\right]\lesssim(n\wedge m)^{-\alpha}\big{(}\log_{+}(n\vee m)\big{)}^{\delta}.$
The hidden constant depends only on $\kappa,\alpha,\epsilon$ and $M$.
Proof.
We follow the notation used in the proof of Theorem 1.1 in [SH25]. In our case, we may take $s=2$, which yields $\beta=\gamma=1/2$. As such, $\alpha\leq\beta$, so the conclusion follows from Comment 1 after the proof of Theorem 1.1 in [SH25]. ∎
A.2. Maximal inequality
The following is taken from Lemma 5 in [CCK15]:
Lemma A.1.
Let $X_{1},\dots,X_{n}$ be independent random vectors in $\mathbb{R}^{k}$ with finite second moments. Set $M=\max_{1\leq i\leq n}\max_{1\leq j\leq k}|X_{ij}|$ and $\sigma^{2}=\max_{1\leq j\leq k}\sum_{i=1}^{n}\mathbb{E}[X_{ij}^{2}]$. Then,
$\mathbb{E}\left[\max_{1\leq j\leq k}\left|\sum_{i=1}^{n}(X_{ij}-\mathbb{E}[X_{ij}])\right|\right]\lesssim\sigma\sqrt{\log_{+}(k)}+\sqrt{\mathbb{E}[M^{2}]}\log_{+}(k)$
up to a universal constant, where $\log_{+}(k)=\log(1+k)$.